# X-Vine Models for Multivariate Extremes

### v1

@@ F.2.1 Box-plots of dependence measure estimates with varying $(n,k)$
F.2.2 Box-plots of pseudo-maximum likelihood estimates of parameters of bivariate tail copulas
Using the same X-vine specification in Fig. 2, we evaluate bias and variance of maximum pseudo-likelihood parameter estimates of tail copula densities. Specifically, we create box-plots of estimates for the Hüsler–Reiss tail copula density associated with the edge $e_{12}\in E_{1}$, using the same combinations of $(n,k)$ defined in Appendix F.2.1. An increase in the effective sample size $n_{D_{e}}$ reduces variability, while a decreasing threshold $k/n$ leads to smaller bias, as expected. Similar results are observed with other tail copula densities (figures are omitted).
@@ F.2.2 Box-plots of pseudo-maximum likelihood estimates of parameters of bivariate tail copulas
F.3 Additional results for US flight delay data in Section 9
F.3.1 Estimated regular vine sequence of the X-vine model from $T_{1}$ to $T_{7}$
@@
Assigning $\widehat{\chi}_{e}$ as edge weight for $e\in E_{1}$ and $\widehat{\tau}_{e}$ for $e\in E_{j}$ with $j\geqslant 2$, we obtain the sequence of maximum spanning trees that satisfy the proximity condition. With the mBIC-optimal truncation level of $q^{*}=7$, Figure 9 displays the trees of the X-vine model from $T_{1}$ to $T_{7}$ sequentially. For trees $T_{2}$ to $T_{7}$, the edges for which the independence copula is selected are not shown.
@@
Figure 9: Sequential display of subsequent trees of the truncated X-vine model. For trees $T_{2}$ to $T_{7}$, the edges for which the independence copula is selected are not shown. (a) $T_{1}$ (b) $T_{2}$ (c) $T_{3}$ (d) $T_{4}$ (e) $T_{5}$ (f) $T_{6}$ (g) $T_{7}$.
F.3.2 Goodness of fit for the fitted (truncated) X-vine model
@@
To assess the goodness of fit for the full X-vine model, we compare empirical pairwise $\widehat{\chi}$ values with those derived from the full X-vine model in Fig. LABEL:fig:ChiflightFullXvine. As an indicator of capturing a higher-dimensional extremal dependence, we use a trivariate (lower) tail dependence coefficient $\chi_{abc}=R_{abc}(1,1,1)$ for distinct $a,b,c\in[d]$. Its empirical estimate is
$\widehat{\chi}_{abc}=\frac{1}{k}\sum_{i=1}^{n}\operatorname{\mathds{1}}\left\{{\widehat{U}_{i,a}\leqslant k/n,\,\widehat{U}_{i,b}\leqslant k/n,\,\widehat{U}_{i,c}\leqslant k/n}\right\}.$
The empirical trivariate tail dependence coefficients are compared with those derived from the full X-vine model in Fig. LABEL:fig:Chi3FlightFull and from the truncated X-vine model in Fig. LABEL:fig:Chi3FlightTrunc, respectively.
@@
Figure 10: A $\chi$-plot comparing empirical pairwise tail dependence coefficients $\widehat{\chi}_{a,b}$ for distinct $a,b\in[d]$ with those derived from the full X-vine model (a), and a plot comparing empirical trivariate tail dependence coefficients $\widehat{\chi}_{a,b,c}$ for distinct $a,b,c\in[d]$ with those obtained from the fitted full X-vine model (b); from the fitted truncated X-vine model with a truncation level of $q^{*}=7$ (c), respectively.
F.3.3 Sensitivity analysis for mBIC truncation levels over quantiles
@@
We conduct a sensitivity analysis to observe how the mBIC-optimal truncation level changes over quantiles. We observe a decline in the mBIC-optimal truncation level as the quantile $k/n$ decreases. The range of mBIC-optimal truncation levels is from 4 to 8 with a median of 6.
@@
Figure 11: (a) mBIC-values plotted across tree levels with a dotted line indicating a selected mBIC-optimal truncation level of $q^{*}=7$. (b) Plot of the mBIC-optimal truncation level across a sequence of quantiles $k/n\in(0.05,0.15)$.
@@
F.3.4 Model comparison: X-vine Hüsler–Reiss model versus Hüsler–Reiss extremal graphical model from EGlearn
@@
Focusing on the Hüsler–Reiss distribution, we compare model assessments between the X-vine Hüsler–Reiss model (all tail copulas in $T_{1}$ are Hüsler–Reiss and all bivariate copulas in the subsequent trees are Gaussian) and the Hüsler–Reiss extremal graphical model obtained with EGlearn algorithm (Engelke et al., 2021). In both models, within $T_{1}$, we use the empirical extremal variogram matrix $\widehat{\Gamma}$ as edge weight for the minimum spanning tree. The resulting tree must be identical for both models, as confirmed in Fig. 13. As described in Section 9, with the tuning parameter value of $\rho^{*}=0.1$ selected by the algorithm, the resulting sparse extremal graph has 148 edges, shown in Fig. LABEL:fig:ChiFlightEGlearn. For the Hüsler–Reiss X-vine, the algorithm sets all bivariate copulas to the independence copula from tree level 19 on (i.e., the truncation level is $q=18$), selecting 217 independence copulas out of 406 edges in trees $T_{1}$ to $T_{28}$ (recall $d=29$). Using the mBIC, we further consider the truncated Hüsler–Reiss X-vine with a lower truncated level. The optimal mBIC truncation level of $q^{*}=7$ results in a total of 175 edges in edges $T_{1}$ to $T_{7}$, shown in Fig. LABEL:fig:mBICFlight. In trees $T_{2}$ to $T_{7}$, there are $46$ edges out of $147$ for which the independence copula is selected.
@@
We assess the goodness of fit between the different graphical models using the entire sub-samples, that is, no longer splitting the data in a training and test set. Fig. 12 compares empirical tail dependence coefficients to those obtained from the fitted models, specifically for the extremal Hüsler–Reiss graphical model based on the flight graph, the truncated Hüsler–Reiss X-vine, and the extremal Hüsler–Reiss graph derived from the EGlearn algorithm, respectively. Similarly to the truncated X-vine model, the truncated Hüsler–Reiss X-vine exhibits improved fit but more variability compared to the extremal Hüsler–Reiss graphical model.
@@
Figure 12: $\chi$-plots comparing empirical pairwise $\widehat{\chi}$ values to those from the fitted graphical models: the extremal Hüsler–Reiss graphical model for the flight graph (a), the truncated Hüsler–Reiss X-vine with $q^{*}=7$ (b), and the extremal Hüsler–Reiss graphical model with $\rho^{*}=0.1$ using the EGlearn algorithm (c), respectively.
F.3.5 Estimated regular vine sequence of the truncated Hüsler–Reiss X-vine
@@
In Fig. 13, we sequentially display the selected trees $T_{1}$ to $T_{7}$ of the truncated Hüsler–Reiss X-vine in Section F.3.4. For trees $T_{2}$ to $T_{7}$, there are $46$ out of $147$ edges for which the independence copula is selected, and these edges are not shown in the figure.
@@
Figure 13: Sequential display of the first $q^{*}=7$ trees of the truncated Hüsler–Reiss X-vine fitted to the US flight data in Section F.3.4. For trees $T_{2}$ to $T_{7}$, the edges for which the independence copula is selected are not shown. (a) $T_{1}$ (b) $T_{2}$ (c) $T_{3}$ (d) $T_{4}$ (e) $T_{5}$ (f) $T_{6}$ (g) $T_{7}$.

### v2

@@ F.1.2 Box-plots of dependence measure estimates with varying $(n,k)$
F.1.3 Box-plots of pseudo-maximum likelihood estimates of parameters of bivariate tail copulas
Using the same X-vine specification in Fig. LABEL:fig:5d_Xvine_a, we evaluate bias and variance of maximum pseudo-likelihood parameter estimates of tail copula densities. Specifically, Fig. F.4 shows box-plots of estimates for the Hüsler–Reiss tail copula density associated with the edge $e_{12}\in E_{1}$, using the same combinations of $(n,k)$ defined in Section F.1.2. An increase in the effective sample size $n_{D_{e}}$ reduces variability, while a decreasing threshold $k/n$ leads to smaller bias, as expected. Similar results are observed with other tail copula densities (figures are omitted).
@@ F.1.3 Box-plots of pseudo-maximum likelihood estimates of parameters of bivariate tail copulas
F.1.4 Averaged-effective sample sizes in different trees
We also investigate how effective sample sizes change over tree levels of Section 8.2. We calculate the averaged-effective sample size over 200 repetitions for each edge in each tree and express it as a percentage of the sample size $n$:
$T_{1}$ 7.30% 6.47% 6.60% 6.87%
$T_{2}$ 5.00% 5.00% 5.00%
$T_{3}$ 3.52% 3.38%
$T_{4}$ 3.12%
The order of the components from left to right and then from top to bottom corresponds to the x-axis order of box-plots in Fig. LABEL:fig:boxplot_Fam. The overall averaged-effective sample size decreases in the first level tree when considering stronger extremal structures (result is omitted). This reduction is due to the fact that, for a high threshold $u$, the probability $\operatorname{\mathsf{P}}[X_{a}>u\text{ or }X_{b}>u]$ decreases as the tail dependence between $X_{a}$ and $X_{b}$ increases. For the second level tree $T_{2}$, the averaged-effective sample is $k/n=5\%$, since the use of the rank transform implies (in the absence of ties) the identity $K_{j}=k$ for all $j\in[d]$.
F.2 Additional results for US flight delay data in Section 9
F.2.1 Estimated regular vine sequence of the X-vine model from $T_{1}$ to $T_{7}$
@@
Assigning $\widehat{\chi}_{e}$ as edge weight for $e\in E_{1}$ and $\widehat{\tau}_{e}$ for $e\in E_{j}$ with $j\geqslant 2$, we obtain the sequence of maximum spanning trees that satisfy the proximity condition. With the mBIC-optimal truncation level of $q^{*}=7$, Figure F.5 displays the trees of the X-vine model from $T_{1}$ to $T_{7}$ sequentially. For trees $T_{2}$ to $T_{7}$, the edges for which the independence copula is selected are not shown.
@@
Figure F.5: Sequential display of subsequent trees of the truncated X-vine model. For trees $T_{2}$ to $T_{7}$, the edges for which the independence copula is selected are not shown. (a) $T_{1}$ (b) $T_{2}$ (c) $T_{3}$ (d) $T_{4}$ (e) $T_{5}$ (f) $T_{6}$ (g) $T_{7}$.
F.2.2 Goodness of fit for the fitted (truncated) X-vine model
@@
To assess the goodness of fit for the full X-vine model, we compare empirical pairwise $\widehat{\chi}$ values with those derived from the full X-vine model in Fig. LABEL:fig:ChiflightFullXvine. As an indicator of capturing a higher-dimensional extremal dependence, we use a trivariate (lower) tail dependence coefficient $\chi_{abc}=R_{abc}(1,1,1)$ for distinct $a,b,c\in[d]$. The empirical trivariate tail dependence coefficients (Section E.2) are compared with those derived from the full X-vine model in Fig. LABEL:fig:Chi3FlightFull and from the truncated X-vine model in Fig. LABEL:fig:Chi3FlightTrunc, respectively.
@@
Figure F.6: (a) A $\chi$-plot comparing empirical pairwise tail dependence coefficients $\widehat{\chi}_{a,b}$ for distinct $a,b\in[d]$ with those derived from the full X-vine model. (b) A plot comparing empirical trivariate tail dependence coefficients $\widehat{\chi}_{a,b,c}$ for distinct $a,b,c\in[d]$ with those obtained from the full X-vine model. (c) similar to (b), but for the truncated X-vine model with a truncation level of $q^{*}=7$.
F.2.3 Sensitivity analysis for mBIC truncation levels over quantiles
@@
Fig. LABEL:fig:mBICFlight illustrates the mBIC-optimal truncation level of $q^{*}=7$. We also conduct a sensitivity analysis to observe how the mBIC-optimal truncation level changes over quantiles. Fig. LABEL:fig:mBICvsQuan shows a decline in the mBIC-optimal truncation level as the quantile $k/n$ decreases. The range of mBIC-optimal truncation levels is from 4 to 8 with a median of 6.
@@
Figure F.7: (a) mBIC-values plotted across tree levels with a dotted line indicating a selected mBIC-optimal truncation level of $q^{*}=7$. (b) Plot of the mBIC-optimal truncation level across a sequence of quantiles $k/n\in(0.05,0.15)$.
@@
F.2.4 Model comparison: X-vine Hüsler–Reiss model versus Hüsler–Reiss extremal graphical model from EGlearn
@@
Focusing on the Hüsler–Reiss distribution, we compare model assessments between the X-vine Hüsler–Reiss model (all tail copulas in $T_{1}$ are Hüsler–Reiss and all bivariate copulas in the subsequent trees are Gaussian) and the Hüsler–Reiss extremal graphical model obtained with EGlearn algorithm (Engelke et al., 2021). In both models, within $T_{1}$, we use the empirical extremal variogram matrix $\widehat{\Gamma}$ as edge weight for the minimum spanning tree. The resulting tree must be identical for both models, as confirmed in Fig. F.9. As described in Section 9, with the tuning parameter value of $\rho^{*}=0.1$ selected by the algorithm, the resulting sparse extremal graph has 148 edges, shown in Fig. LABEL:fig:ChiFlightEGlearn. For the Hüsler–Reiss X-vine, the algorithm sets all bivariate copulas to the independence copula from tree level 19 on (i.e., the truncation level is $q=18$), selecting 217 independence copulas out of 406 edges in trees $T_{1}$ to $T_{28}$ (recall $d=29$). Using the mBIC, we further consider the truncated Hüsler–Reiss X-vine with a lower truncated level. The optimal mBIC truncation level of $q^{*}=7$ results in a total of 175 edges in edges $T_{1}$ to $T_{7}$, shown in Fig. LABEL:fig:mBICFlight. In trees $T_{2}$ to $T_{7}$, there are $46$ edges out of $147$ for which the independence copula is selected.
We assess the goodness of fit between the different graphical models using the entire sub-sample, that is, we no longer split the data in a training and test set. As in Fig. 7 in Hentschel et al. (2022) and Fig. 8(e) in Engelke et al. (2024), an alternative approach to creating a $\chi$-plot for the Hüsler–Reiss distribution is to convert the variogram elements $\Gamma_{ab}$ for distinct $a,b\in[d]$ to the tail dependence coefficients using the relationship $\chi_{ab}=2-2\Phi(\sqrt{\Gamma_{ab}}/2)$ for distinct $a,b\in[d]$.
@@
Fig. F.8 compares the tail dependence coefficients converted from the empirical variogram to those from the fitted variogram. This comparision is done specifically for the extremal Hüsler–Reiss graphical model based on the flight graph, the truncated Hüsler–Reiss X-vine, and the extremal Hüsler–Reiss graph derived from the EGlearn algorithm, respectively. For the truncated Hüsler–Reiss X-vine, a similar procedure to steps 1 and 2 in Section E.2.3 is employed to obtain the fitted variogram matrix and subsequently transform it into fitted $\chi$ values. Similarly, like the truncated X-vine model, the truncated Hüsler–Reiss X-vine exhibits more variability compared to the extremal Hüsler–Reiss graphical model.
@@
Figure F.8: $\chi$-plots comparing pairwise $\widehat{\chi}$ values converted from the empirical variogram to those from the fitted variogram for graphical models: (a) the extremal Hüsler–Reiss graphical model for the flight graph, (b) the truncated Hüsler–Reiss X-vine with $q^{*}=7$, and (c) the extremal Hüsler–Reiss graphical model with $\rho^{*}=0.1$ using the EGlearn algorithm.
F.2.5 Estimated regular vine sequence of the truncated Hüsler–Reiss X-vine
@@
In Fig. F.9, we sequentially display the selected trees $T_{1}$ to $T_{7}$ of the truncated Hüsler–Reiss X-vine in Section F.2.4. For trees $T_{2}$ to $T_{7}$, there are $46$ out of $147$ edges for which the independence copula is selected, and these edges are not shown in the figure.
@@
Figure F.9: Sequential display of the first $q^{*}=7$ trees of the truncated Hüsler–Reiss X-vine fitted to the US flight data in Section F.2.4. For trees $T_{2}$ to $T_{7}$, the edges for which the independence copula is selected are not shown. (a) $T_{1}$ (b) $T_{2}$ (c) $T_{3}$ (d) $T_{4}$ (e) $T_{5}$ (f) $T_{6}$ (g) $T_{7}$.